# Contraherent cosheaves of contramodules on Noetherian formal schemes

### 29 Mar 2026 v1

@@ 2.13. Colocality of exactness
Part (b): the “only if” assertion follow from part (a). To prove the “if”, notice that the map $\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{N})\longrightarrow\mathfrak{N}$ is an admissible epimorphism in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ by Lemma 2.12.11. Here we are using the fact that the full subcategory $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ is closed under cokernels (in fact, even under quotients) in $R{\operatorname{\mathsf{--Contra}}}$. If the map $f_{\alpha}^{!}(\mathfrak{M})\longrightarrow f_{\alpha}^{!}(\mathfrak{N})$ is an admissible epimorphism in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$, then the map $f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{M})\longrightarrow f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{N})$ is an admissible epimorphism in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ by Lemma 2.9.7. Hence the composition $\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{M})\longrightarrow\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{N})\longrightarrow\mathfrak{N}$ is an admissible epimorphism in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$. The latter composition is equal to the composition $\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{M})\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}$, and it follows that $\mathfrak{M}\longrightarrow\mathfrak{N}$ is an admissible epimorphism in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ by [4, Proposition 7.6].
Part (c): the “only if” assertion follow from part (a). To prove the “if”, notice that if the complexes $f_{\alpha}^{!}(\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ are exact in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$, then the complexes $f_{\alpha_{1},\dotsc,\alpha_{k}}^{!}(\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ are exact in $S_{\alpha_{1},\dotsc,\alpha_{k}}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ for all $k\geq 1$ (by the “only if” assertion). Hence the complexes $f_{\alpha_{1},\dotsc,\alpha_{k}}{}_{\sharp}f_{\alpha_{1},\dotsc,\alpha_{k}}^{!}(\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ are exact in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ (by Lemma 2.9.7). Applying the construction of the finite exact sequence (10) from Lemma 2.12.11 to every term of the complex $\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, we obtain a finite termwise exact sequence of complexes in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$. Since all the complexes appearing in the latter finite termwise exact sequence, except perhaps the rightmost one, are exact in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$, it follows that the rightmost complex $\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$, too (e. g., by [4, Corollary 3.6(i)]). ∎
@@
Notice that the obvious analogues of Lemmas 2.13.9(b), 2.13.10(b), and 2.13.11(b) for admissible (or split) monomorphisms are not true (not even for discrete regular Noetherian commutative rings $R$ and $S_{\alpha}$ of Krull dimension $1$). The counterexample can be discerned from [17, Example 3.2.1].
@@
Part (a) is easy: the assertion follows from the fact that the full subcategory $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ is closed under quotients in $R{\operatorname{\mathsf{--Contra}}}$. Part (b) is the cotorsion periodicity theorem for contramodules; see [27, Corollary 12.4]. ∎
@@ 3.1. The basics
Let $\mathfrak{X}$ be a locally Noetherian formal scheme and $\mathfrak{O}=\mathfrak{O}_{\mathfrak{X}}$ be its structure sheaf. For every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$, the ring of functions $\mathfrak{R}=\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$ is a Noetherian commutative ring endowed with a complete, separated adic topology (in the sense of Section 2.1).
@@
Closed subschemes of definition $\overline{X}\subset\mathfrak{X}$ correspond bijectively to coherent sheaves of ideals of definition. The latter term means coherent sheaves of ideals $\mathfrak{I}$ on $\mathfrak{X}$ such that, for every affine open subscheme $\mathfrak{U}\subset\mathfrak{X}$, the ideal $\mathfrak{I}(\mathfrak{U})\subset\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$ is an ideal of definition in the adic Noetherian ring $\mathfrak{R}=\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$. In particular, the minimal closed subscheme of definition $\overline{X}_{\mathrm{mn}}\subset\mathfrak{X}$ corresponds to the quasi-coherent sheaf formed by the maximal ideals of definition $\mathfrak{I}_{\mathrm{mx}}\subset\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$.
@@
Given an open covering $\mathbf{W}$ of a locally Noetherian formal scheme $\mathfrak{X}$, an open formal subscheme $\mathfrak{Y}\subset\mathfrak{X}$ is said to be subordinate to $\mathbf{W}$ if there exists an open subscheme $\mathfrak{W}\in\mathbf{W}$ for which $\mathfrak{Y}\subset\mathfrak{W}$. An open covering $\mathfrak{X}=\bigcup_{\alpha}\mathfrak{Y}_{\alpha}$ is said to be subordinate to $\mathbf{W}$ if $\mathfrak{Y}_{\alpha}$ is subordinate to $\mathbf{W}$ for every $\alpha$. More generally, the same terminology is applied to open coverings of arbitrary topological spaces.
@@ 3.2. Cosheaves of modules over a ringed space
Let $(\mathfrak{X},\mathfrak{O})$ be a ringed space and $\mathbf{B}$ be a base of open subsets in $\mathfrak{X}$. Recall the definitions of a copresheaf of $\mathfrak{O}$-modules on $\mathbf{B}$ and a cosheaf of $\mathfrak{O}$-modules on $\mathbf{B}$ from [17, Section 2.1] and [29, Section 2.1].
@@
The main result in connection with the definition of a sheaf of $\mathfrak{O}$-modules on $\mathfrak{X}$ is the following theorem from [11, Section 0.3.2], [13, Lemma Tag 009U], [17, Proposition 2.1.3], [29, Theorem 2.1(a)]: the natural restriction functor $(\mathfrak{X},\mathfrak{O}){\operatorname{\mathsf{--Sh}}}\longrightarrow(\mathbf{B},\mathfrak{O}){\operatorname{\mathsf{--Sh}}}$ is a category equivalence. So any sheaf of $\mathfrak{O}$-modules on $\mathbf{B}$ can be extended to a sheaf of $\mathfrak{O}$-modules on $\mathfrak{X}$ in a unique way.
@@
The main result in connection with the definition of a cosheaf of $\mathfrak{O}$-modules on $\mathfrak{X}$ is the following theorem from [17, Theorem 2.1.2], [29, Theorem 2.1(b)]: the natural restriction functor $(\mathfrak{X},\mathfrak{O}){\operatorname{\mathsf{--Cosh}}}\longrightarrow(\mathbf{B},\mathfrak{O}){\operatorname{\mathsf{--Cosh}}}$ is a category equivalence. So any cosheaf of $\mathfrak{O}$-modules on $\mathbf{B}$ can be extended to a cosheaf of $\mathfrak{O}$-modules on $\mathfrak{X}$ in a unique way.
@@
(2) Assume that the base of open subsets $\mathbf{B}$ in $\mathfrak{X}$ consists of quasi-compact open subsets, and furthermore, the intersection of any two open subsets from $\mathbf{B}$ that are contained in a third open subset from $\mathbf{B}$ is also quasi-compact. Then, following [17, Remark 2.1.4], it suffices to check axioms (11) and (12) for finite open coverings $\mathfrak{U}=\bigcup_{\alpha}\mathfrak{V}_{\alpha}$ and finite open coverings of the intersections $\mathfrak{V}_{\alpha}\cap\mathfrak{V}_{\beta}=\bigcup_{\gamma}\mathfrak{W}_{\alpha\beta\gamma}$.
@@ 3.3. Cosheaves of contramodules over a formal scheme
For any Noetherian commutative ring $R$ with an adic topology, we denote by $R{\operatorname{\mathsf{--Tors}}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\subset R{\operatorname{\mathsf{--Mod}}}$ the full subcategory of torsion $R$-modules and by $R{\operatorname{\mathsf{--Contra}}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\subset R{\operatorname{\mathsf{--Mod}}}$ the full subcategory of contramodule $R$-modules. All contramodule $R$-modules are quotseparated due to the assumption that the ring $R$ is Noetherian. See the notation and discussion in Sections 2.2, 2.4, and 2.9.
@@
Let $\mathfrak{X}$ be a locally Noetherian formal scheme. A sheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules $\mathcal{M}$ on $\mathfrak{X}$ is said to be a sheaf of torsion $\mathfrak{O}_{\mathfrak{X}}$-modules if, for every affine open subscheme $\mathfrak{U}\subset\mathfrak{X}$, the $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-module $\mathcal{M}(\mathfrak{U})$ is torsion/discrete as a module over a topological commutative ring with an adic topology. This means that the annihilator of any element of $\mathcal{M}(\mathfrak{U})$ is an open ideal in $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$, or equivalently, for every ideal of definition $\mathfrak{I}\subset\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$ and every element $x\in\mathcal{M}(\mathfrak{U})$ there exists an integer $n\geq 1$ such that $\mathfrak{I}^{n}x=0$ in $\mathcal{M}(\mathfrak{U})$.
@@
The sheaf axiom for $\mathcal{M}$ implies that it suffices to check the torsion condition for affine open subschemes $\mathfrak{U}\subset\mathfrak{X}$ subordinate to a given open covering of $\mathfrak{X}$. Indeed, if $\mathfrak{U}=\bigcup_{\alpha}\mathfrak{V}_{\alpha}$ is a finite affine open covering of an affine formal scheme $\mathfrak{U}$ and $\mathcal{N}$ is a sheaf of $\mathfrak{O}_{\mathfrak{U}}$-modules on $\mathfrak{U}$, then $\mathcal{N}(\mathfrak{U})$ is a torsion module over $\mathfrak{R}=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$ whenever $\mathcal{N}(\mathfrak{V}_{\alpha})$ is a torsion module over $\mathfrak{S}_{\alpha}=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}_{\alpha})$ for all indices $\alpha$. The latter assertion follows from the facts that the map $\mathcal{N}(\mathfrak{U})\longrightarrow\bigoplus_{\alpha}\mathcal{N}(\mathfrak{V}_{\alpha})$ is injective, the functors of restrictions of scalars $f_{\alpha}{}_{*}\colon\mathfrak{S}_{\alpha}{\operatorname{\mathsf{--Mod}}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}$ take $\mathfrak{S}_{\alpha}{\operatorname{\mathsf{--Tors}}}$ into $\mathfrak{R}{\operatorname{\mathsf{--Tors}}}$ (see Section 2.6), and any submodule of a torsion $\mathfrak{R}$-module is torsion.
@@
A cosheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules $\mathfrak{P}$ on $\mathfrak{X}$ is said to be a cosheaf of contramodule $\mathfrak{O}_{\mathfrak{X}}$-modules if, for every affine open subscheme $\mathfrak{U}\subset\mathfrak{X}$, the $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-module $\mathfrak{P}[\mathfrak{U}]$ is a contramodule over the adic topological ring $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$. We refer to Section 2.4 for the definition of contramodule $R$-modules for an adic topological ring $R$. Let us only point out that, since the commutative ring $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$ is Noetherian by assumption, all the contramodule $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules are quotseparated.
@@
The cosheaf axiom for $\mathfrak{P}$ implies that is suffices to check the contramoduleness condition for affine open subschemes $\mathfrak{U}\subset\mathfrak{X}$ subordinate to a given open covering of $\mathfrak{X}$. Indeed, let $\mathfrak{U}=\bigcup_{\alpha}\mathfrak{V}_{\alpha}$ be a finite affine open covering of an affine formal scheme $\mathfrak{U}$. Let $\mathfrak{Q}$ be a cosheaf of $\mathfrak{O}_{\mathfrak{U}}$-modules on $\mathfrak{U}$. Then the $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$-module $\mathfrak{Q}[\mathfrak{U}]$ is a contramodule over the adic topological ring $\mathfrak{R}=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$ whenever the $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}_{\alpha})$-module $\mathfrak{Q}[\mathfrak{V}_{\alpha}]$ is a contramodule over the adic topological ring $\mathfrak{S}_{\alpha}=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}_{\alpha})$ and the $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}_{\alpha,\beta})$-module $\mathfrak{Q}[\mathfrak{V}_{\alpha,\beta}]$ is a contramodule over the adic topological ring $\mathfrak{S}_{\alpha,\beta}=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}_{\alpha,\beta})$ for all indices $\alpha$ and $\beta$. Here we put $\mathfrak{V}_{\alpha,\beta}=\mathfrak{V}_{\alpha}\cap\mathfrak{V}_{\beta}\subset\mathfrak{U}$.

### 17 Apr 2026 v2

@@ 2.17. Colocality of exactness
Part (b): the “only if” assertion follow from part (a). To prove the “if”, notice that the map $\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{N})\longrightarrow\mathfrak{N}$ is an admissible epimorphism in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ by Lemma 2.16.11. Here we are using the fact that the full subcategory $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ is closed under cokernels (in fact, even under quotients) in $R{\operatorname{\mathsf{--Contra}}}$. If the map $f_{\alpha}^{!}(\mathfrak{M})\longrightarrow f_{\alpha}^{!}(\mathfrak{N})$ is an admissible epimorphism in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$, then the map $f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{M})\longrightarrow f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{N})$ is an admissible epimorphism in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ by Lemma 2.12.7. Hence the composition $\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{M})\longrightarrow\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{N})\longrightarrow\mathfrak{N}$ is an admissible epimorphism in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$. The latter composition is equal to the composition $\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{!}(\mathfrak{M})\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}$, and it follows that $\mathfrak{M}\longrightarrow\mathfrak{N}$ is an admissible epimorphism in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ by [5, Proposition 7.6].
Part (c): the “only if” assertion follow from part (a). To prove the “if”, notice that if the complexes $f_{\alpha}^{!}(\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ are exact in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$, then the complexes $f_{\alpha_{1},\dotsc,\alpha_{k}}^{!}(\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ are exact in $S_{\alpha_{1},\dotsc,\alpha_{k}}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ for all $k\geq 1$ (by the “only if” assertion). Hence the complexes $f_{\alpha_{1},\dotsc,\alpha_{k}}{}_{\sharp}f_{\alpha_{1},\dotsc,\alpha_{k}}^{!}(\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ are exact in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ (by Lemma 2.12.7). Applying the construction of the finite exact sequence (10) from Lemma 2.16.11 to every term of the complex $\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, we obtain a finite termwise exact sequence of complexes in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$. Since all the complexes appearing in the latter finite termwise exact sequence, except perhaps the rightmost one, are exact in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$, it follows that the rightmost complex $\mathfrak{M}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is exact in $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$, too (e. g., by [5, Corollary 3.6(i)]). ∎
@@
Notice that the obvious analogues of Lemmas 2.17.9(b), 2.17.10(b), and 2.17.11(b) for admissible (or split) monomorphisms are not true (not even for discrete regular Noetherian commutative rings $R$ and $S_{\alpha}$ of Krull dimension $1$). The counterexample can be discerned from [18, Example 3.2.1].
@@
Part (a) is easy: the assertion follows from the fact that the full subcategory $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ is closed under quotients in $R{\operatorname{\mathsf{--Contra}}}$. Part (b) is the cotorsion periodicity theorem for contramodules; see [28, Corollary 12.4]. ∎
@@ 3.1. The basics
Let $\mathfrak{X}$ be a locally Noetherian formal scheme and $\mathfrak{O}=\mathfrak{O}_{\mathfrak{X}}$ be its structure sheaf. For every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$, the ring of functions $\mathfrak{R}=\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$ is a Noetherian commutative ring endowed with a complete, separated adic topology (in the sense of Section 2.1). The results collected in Section 2.9 are relevant to the fact that the notion of a (locally) Noetherian formal scheme is reasonably well-behaved.
@@
Closed subschemes of definition $\overline{X}\subset\mathfrak{X}$ correspond bijectively to coherent sheaves of ideals of definition. The latter term means coherent sheaves of ideals $\mathfrak{I}$ on $\mathfrak{X}$ such that, for every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$, the ideal $\mathfrak{I}(\mathfrak{U})\subset\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$ is an ideal of definition in the adic Noetherian ring $\mathfrak{R}=\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$. In particular, the minimal closed subscheme of definition $\overline{X}_{\mathrm{mn}}\subset\mathfrak{X}$ corresponds to the quasi-coherent sheaf formed by the maximal ideals of definition $\mathfrak{I}_{\mathrm{mx}}\subset\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$.
@@
A locally Noetherian formal scheme $\mathfrak{X}$ is said to be Noetherian if $\mathfrak{X}$ is quasi-compact (i. e., the underlying topological space of $\mathfrak{X}$ is quasi-compact). A locally Noetherian formal scheme $\mathfrak{X}$ is said to be semi-separated if the intersection of any two affine open formal subschemes in $\mathfrak{X}$ is an affine open formal subscheme. Notice that every locally Noetherian formal subschemes $\mathfrak{X}$ is quasi-separated in the sense that the intersection of any two quasi-compact open formal subschemes in $\mathfrak{X}$ is quasi-compact.
Let $\mathfrak{X}$ be a locally Noetherian formal scheme and $\overline{X}\subset\mathfrak{X}$ be a closed subscheme of definition in $\mathfrak{X}$. Then the formal scheme $\mathfrak{X}$ is affine if and only if the scheme $\overline{X}$ is affine. Indeed, the underlying topological spaces of $\mathfrak{X}$ and $\overline{X}$ coincide; hence without loss of generality we can assume $\mathfrak{X}$ and $\overline{X}$ to be quasi-compact (i. e., Noetherian). Then the assertion follows from the fact that, for any Noetherian scheme $X$ with a closed subscheme $\overline{X}$, the scheme $X$ is affine if and only if $\overline{X}$ is affine [13, Exercise III.3.1], [14, Lemma Tag 01YQ].
Consequently, a locally Noetherian formal scheme $\mathfrak{X}$ is semi-separated if and only if it has a semi-separated closed subscheme of definition $\overline{X}$, or equivalently, if and only if every closed subscheme of definition in $\mathfrak{X}$ is semi-separated.
Given an open covering $\mathbf{W}$ of a locally Noetherian formal scheme $\mathfrak{X}$, an open formal subscheme $\mathfrak{Y}\subset\mathfrak{X}$ is said to be subordinate to $\mathbf{W}$ if there exists an open formal subscheme $\mathfrak{W}\in\mathbf{W}$ for which $\mathfrak{Y}\subset\mathfrak{W}$. An open covering $\mathfrak{X}=\bigcup_{\alpha}\mathfrak{Y}_{\alpha}$ is said to be subordinate to $\mathbf{W}$ if $\mathfrak{Y}_{\alpha}$ is subordinate to $\mathbf{W}$ for every $\alpha$. More generally, the same terminology is applied to open coverings of arbitrary topological spaces.
@@ 3.2. Cosheaves of modules over a ringed space
Let $(\mathfrak{X},\mathfrak{O})$ be a ringed space and $\mathbf{B}$ be a base of open subsets in $\mathfrak{X}$. Recall the definitions of a copresheaf of $\mathfrak{O}$-modules on $\mathbf{B}$ and a cosheaf of $\mathfrak{O}$-modules on $\mathbf{B}$ from [18, Section 2.1] and [30, Section 2.1].
@@
The main result in connection with the definition of a sheaf of $\mathfrak{O}$-modules on $\mathfrak{X}$ is the following theorem from [12, Section 0.3.2], [14, Lemma Tag 009U], [18, Proposition 2.1.3], [30, Theorem 2.1(a)]: the natural restriction functor $(\mathfrak{X},\mathfrak{O}){\operatorname{\mathsf{--Sh}}}\longrightarrow(\mathbf{B},\mathfrak{O}){\operatorname{\mathsf{--Sh}}}$ is a category equivalence. So any sheaf of $\mathfrak{O}$-modules on $\mathbf{B}$ can be extended to a sheaf of $\mathfrak{O}$-modules on $\mathfrak{X}$ in a unique way.
@@
The main result in connection with the definition of a cosheaf of $\mathfrak{O}$-modules on $\mathfrak{X}$ is the following theorem from [18, Theorem 2.1.2], [30, Theorem 2.1(b)]: the natural restriction functor $(\mathfrak{X},\mathfrak{O}){\operatorname{\mathsf{--Cosh}}}\longrightarrow(\mathbf{B},\mathfrak{O}){\operatorname{\mathsf{--Cosh}}}$ is a category equivalence. So any cosheaf of $\mathfrak{O}$-modules on $\mathbf{B}$ can be extended to a cosheaf of $\mathfrak{O}$-modules on $\mathfrak{X}$ in a unique way.
@@
(2) Assume that the base of open subsets $\mathbf{B}$ in $\mathfrak{X}$ consists of quasi-compact open subsets, and furthermore, the intersection of any two open subsets from $\mathbf{B}$ that are contained in a third open subset from $\mathbf{B}$ is also quasi-compact. Then, following [18, Remark 2.1.4], it suffices to check axioms (11) and (12) for finite open coverings $\mathfrak{U}=\bigcup_{\alpha}\mathfrak{V}_{\alpha}$ and finite open coverings of the intersections $\mathfrak{V}_{\alpha}\cap\mathfrak{V}_{\beta}=\bigcup_{\gamma}\mathfrak{W}_{\alpha\beta\gamma}$.
@@ 3.3. Cosheaves of contramodules over a formal scheme
For any Noetherian commutative ring $R$ with an adic topology, we denote by $R{\operatorname{\mathsf{--Tors}}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{tors}}\subset R{\operatorname{\mathsf{--Mod}}}$ the full subcategory of torsion $R$-modules and by $R{\operatorname{\mathsf{--Contra}}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}^{\mathsf{qs}}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{ctra}}\subset R{\operatorname{\mathsf{--Mod}}}$ the full subcategory of contramodule $R$-modules. All contramodule $R$-modules are quotseparated due to the assumption that the ring $R$ is Noetherian. See the notation and discussion in Sections 2.2, 2.4, and 2.12.
@@
Let $\mathfrak{X}$ be a locally Noetherian formal scheme. A sheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules $\mathcal{M}$ on $\mathfrak{X}$ is said to be a sheaf of torsion $\mathfrak{O}_{\mathfrak{X}}$-modules if, for every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$, the $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-module $\mathcal{M}(\mathfrak{U})$ is torsion/discrete as a module over a topological commutative ring with an adic topology. This means that the annihilator of any element of $\mathcal{M}(\mathfrak{U})$ is an open ideal in $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$, or equivalently, for every ideal of definition $\mathfrak{I}\subset\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$ and every element $x\in\mathcal{M}(\mathfrak{U})$ there exists an integer $n\geq 1$ such that $\mathfrak{I}^{n}x=0$ in $\mathcal{M}(\mathfrak{U})$.
@@
The sheaf axiom for $\mathcal{M}$ implies that it suffices to check the torsion condition for affine open formal subschemes $\mathfrak{U}\subset\mathfrak{X}$ subordinate to a given open covering of $\mathfrak{X}$. Indeed, if $\mathfrak{U}=\bigcup_{\alpha}\mathfrak{V}_{\alpha}$ is a finite affine open covering of an affine formal scheme $\mathfrak{U}$ and $\mathcal{N}$ is a sheaf of $\mathfrak{O}_{\mathfrak{U}}$-modules on $\mathfrak{U}$, then $\mathcal{N}(\mathfrak{U})$ is a torsion module over $\mathfrak{R}=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$ whenever $\mathcal{N}(\mathfrak{V}_{\alpha})$ is a torsion module over $\mathfrak{S}_{\alpha}=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}_{\alpha})$ for all indices $\alpha$. The latter assertion follows from the facts that the map $\mathcal{N}(\mathfrak{U})\longrightarrow\bigoplus_{\alpha}\mathcal{N}(\mathfrak{V}_{\alpha})$ is injective, the functors of restrictions of scalars $f_{\alpha}{}_{*}\colon\mathfrak{S}_{\alpha}{\operatorname{\mathsf{--Mod}}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Mod}}}$ take $\mathfrak{S}_{\alpha}{\operatorname{\mathsf{--Tors}}}$ into $\mathfrak{R}{\operatorname{\mathsf{--Tors}}}$ (see Section 2.6), and any submodule of a torsion $\mathfrak{R}$-module is torsion.
@@
A cosheaf of $\mathfrak{O}_{\mathfrak{X}}$-modules $\mathfrak{P}$ on $\mathfrak{X}$ is said to be a cosheaf of contramodule $\mathfrak{O}_{\mathfrak{X}}$-modules if, for every affine open formal subscheme $\mathfrak{U}\subset\mathfrak{X}$, the $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-module $\mathfrak{P}[\mathfrak{U}]$ is a contramodule over the adic topological ring $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$. We refer to Section 2.4 for the definition of contramodule $R$-modules for an adic topological ring $R$. Let us only point out that, since the commutative ring $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$ is Noetherian by assumption, all the contramodule $\mathfrak{O}_{\mathfrak{X}}(\mathfrak{U})$-modules are quotseparated.
@@
The cosheaf axiom for $\mathfrak{P}$ implies that is suffices to check the contramoduleness condition for affine open formal subschemes $\mathfrak{U}\subset\mathfrak{X}$ subordinate to a given open covering of $\mathfrak{X}$. Indeed, let $\mathfrak{U}=\bigcup_{\alpha}\mathfrak{V}_{\alpha}$ be a finite affine open covering of an affine formal scheme $\mathfrak{U}$. Let $\mathfrak{Q}$ be a cosheaf of $\mathfrak{O}_{\mathfrak{U}}$-modules on $\mathfrak{U}$. Then the $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$-module $\mathfrak{Q}[\mathfrak{U}]$ is a contramodule over the adic topological ring $\mathfrak{R}=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{U})$ whenever the $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}_{\alpha})$-module $\mathfrak{Q}[\mathfrak{V}_{\alpha}]$ is a contramodule over the adic topological ring $\mathfrak{S}_{\alpha}=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}_{\alpha})$ and the $\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}_{\alpha,\beta})$-module $\mathfrak{Q}[\mathfrak{V}_{\alpha,\beta}]$ is a contramodule over the adic topological ring $\mathfrak{S}_{\alpha,\beta}=\mathfrak{O}_{\mathfrak{U}}(\mathfrak{V}_{\alpha,\beta})$ for all indices $\alpha$ and $\beta$. Here we put $\mathfrak{V}_{\alpha,\beta}=\mathfrak{V}_{\alpha}\cap\mathfrak{V}_{\beta}\subset\mathfrak{U}$.